### 01 Feb 2024 v1

@@ Appendix A Msss matyices
$\Delta_{(2+i)6}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (111)
×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
@@
$\Delta_{(2+i)7}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})$ (112)
×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}},$
@@
$\Delta_{(2+i)(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})$ (113)
×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})-\Re(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}},$
@@
$\Delta_{16}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}$ (114)
@@
×$\times\Bigg{\{}\ln{m_{\tilde{b}_{1}}^{2}\over m_{\tilde{b}_{2}}^{2}}+{{|A_{b}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})\Bigg{\}},$
@@
$\Delta_{17}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (115)
@@
×$\times\Bigg{\{}\ln{m_{\tilde{b}_{1}}^{2}\over m_{\tilde{b}_{2}}^{2}}+{{|A_{b}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}},$
@@
$\Delta_{26}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}$ (116)
×$\times{{|\mu|^{2}\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$
@@
$\Delta_{27}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (117)
×$\times{{|\mu|^{2}\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}},$
@@
$\Delta_{1(7+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (118)
@@
×$\times\Bigg{\{}\ln{m_{\tilde{b}_{1}}^{2}\over m_{\tilde{b}_{2}}^{2}}+{{|A_{b}|^{2}-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}}$
@@
$\Delta_{2(7+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (119)
×$\times{{|\mu|^{2}\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}},$
@@
$\Delta_{(2+i)6}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (120)
×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}},$
@@
$\Delta_{(2+i)7}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (121)
×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}},$
@@
$\Delta_{(2+i)(7+j)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\Im(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (122)
×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\Re(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}.$

### 06 Mar 2024 v2

@@ Acknowledgements.
$\Delta_{(2+i)6}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}$ (111)
×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2}),$
@@
$\Delta_{(2+i)7}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\mu\cot\beta}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})$ (112)
×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}},$
@@
$\Delta_{(2+i)(7+i)}^{t}\!\!\!\!\!$ $={{3G_{F}m_{t}^{4}}\over{2\sqrt{2}\pi^{2}\sin^{2}\beta}}{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}}g(m_{\tilde{t}_{1}}^{2},m_{\tilde{t}_{2}}^{2})$ (113)
×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{d}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{A_{t}}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{t}\lambda_{i}v_{d}}\over{m_{\tilde{t}_{1}}^{2}-m_{\tilde{t}_{2}}^{2}}},$
@@
$\Delta_{16}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}$ (114)
@@
×$\times\Bigg{\{}\ln{m_{\tilde{b}_{1}}^{2}\over m_{\tilde{b}_{2}}^{2}}+{{|A_{b}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})\Bigg{\}},$
@@
$\Delta_{17}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (115)
@@
×$\times\Bigg{\{}\ln{m_{\tilde{b}_{1}}^{2}\over m_{\tilde{b}_{2}}^{2}}+{{|A_{b}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}},$
@@
$\Delta_{26}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}$ (116)
×$\times{{|\mu|^{2}\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2}),$
@@
$\Delta_{27}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (117)
×$\times{{|\mu|^{2}\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}},$
@@
$\Delta_{1(7+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (118)
@@
×$\times\Bigg{\{}\ln{m_{\tilde{b}_{1}}^{2}\over m_{\tilde{b}_{2}}^{2}}+{{|A_{b}|^{2}-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}\Bigg{\}}$
@@
$\Delta_{2(7+i)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (119)
×$\times{{|\mu|^{2}\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}},$
@@
$\Delta_{(2+i)6}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{-\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu\tan\beta}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (120)
×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}},$
@@
$\Delta_{(2+i)7}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\mu}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (121)
×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}},$
@@
$\Delta_{(2+i)(7+j)}^{b}\!\!\!\!\!$ $={{3G_{F}m_{b}^{4}}\over{2\sqrt{2}\pi^{2}\cos^{2}\beta}}{{\text{Im}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}g(m_{\tilde{b}_{1}}^{2},m_{\tilde{b}_{2}}^{2})$ (122)
×$\times{{{1\over 2}\upsilon_{\nu_{i}^{c}}v_{u}(\lambda_{i}\lambda_{j}^{*}+\lambda_{i}^{*}\lambda_{j})\tan\beta-\text{Re}(e^{i\phi_{\lambda}}e^{i\phi_{v_{u}}}e^{-i\phi_{v_{d}}}e^{i\phi_{\upsilon_{\nu_{i}^{c}}}})A_{b}\lambda_{i}v_{u}}\over{m_{\tilde{b}_{1}}^{2}-m_{\tilde{b}_{2}}^{2}}}.$